OpenStudy (anonymous):

what is the standard form equation of the line passing though (3,-2) with a slope of 3/5

OpenStudy (katrinakaif):

Make a standard equation by using y =mx + b M is the slope which you already have : 3/5 Plug the ordered pair into this equation as (3, - 2) = (x1, y1) (y - y1) = m (x - x1)